OpenStudy (anonymous):

is anyone familiar with the unit of force: kilogramweight? how is it related to newtons

OpenStudy (ganpat):

F = Ma M = mass = Kg a = acceleration = m/s2 therefore, F = Kg * m/s2 = N N = newton

OpenStudy (shubhamsrg):

weight is just another force..which we apply on ground .. and thus is the definiton..

OpenStudy (anonymous):

yup 1N=1kgms-2

OpenStudy (ganpat):

nope....

OpenStudy (anonymous):

but whats kilogram weight?? 1kgweight=gN 1N=1/g kg weight

OpenStudy (anonymous):

nope?

OpenStudy (ganpat):

whats acceleration ?? i mean what is gravitational force of the earth ?

OpenStudy (anonymous):

? what do u mean

OpenStudy (ganpat):

g = gravitational force... what is that ???

OpenStudy (anonymous):

gravitational constant 9.8m/s2?

OpenStudy (ganpat):

right...

OpenStudy (ganpat):

F = Ma assume, F = 1N and M = 1Kg Therefore, 1 N = 9.81 Kg/ m2

OpenStudy (ganpat):

understood ??

OpenStudy (anonymous):

oh so you just use f=ma F=1N, a=g and m=1kg

OpenStudy (ganpat):

yup...

OpenStudy (anonymous):

wait so.. how to get 1N=1/g kgweight?

OpenStudy (ganpat):

ok..

OpenStudy (ganpat):

u sure, its 1/ g ??

OpenStudy (anonymous):

Units: SI unit: Newton (N). A net force of 1 N causes a mass of 1 kg to accelerate at 1 ms-2. Alternative unit: Kilogram weight (kg wt). A force of 1 kg wt is the force exerted by gravity on a mass of 1 kg. 1 Conversion of units: 1 kg wt = g N, 1 N = 1/g kg wt .

OpenStudy (anonymous):

idunno thats what it says

OpenStudy (shubhamsrg):

it is simply equal to the magnitude of the force exerted by one kg of mass in a 9.8 ms^-2 gravitational field(on earth) 1KgF = 9.80665 N and F and weight are congruent..

OpenStudy (ganpat):

lil confusing for me as well...